# LYT-NET: LIGHTWEIGHT YUV TRANSFORMER-BASED NETWORK FOR LOW-LIGHT IMAGE ENHANCEMENT

### v1

@@ 1 Introduction
In this paper, we propose a novel transformer-based approach, characterized by its low-weight design, that achieves state-of-the-art (SOTA) results in LLIE while maintaining computational efficiency. In Fig. 1 we present a comparative analysis of performance over complexity between SOTA methods evaluated using the LOL dataset [2]. For visual clarity, in the short comparison, we omitted models such as Restormer[12] and MIRNet[13] as they utilize considerably more parameters yet yield inferior results.
@@ 3.3 Channel-wise Denoiser Block
FLOPS (G) Params (M) PSNR SSIM PSNR SSIM PSNR SSIM
  SID [23] 13.73 7.76 14.35 0.436 13.24 0.442 15.04 0.610
3DLUT [24] 0.075 0.59 14.35 0.445 17.59 0.721 18.04 0.800
DeepUPE [25] 21.10 1.02 14.38 0.446 13.27 0.452 15.08 0.623
DeepLPF [26] 5.86 1.77 15.28 0.473 14.10 0.480 16.02 0.587
UFormer [27] 12.00 5.29 16.36 0.771 18.82 0.771 19.66 0.871
RetinexNet [2] 587.47 0.84 16.77 0.560 15.47 0.567 17.13 0.798
Sparse [28] 53.26 2.33 17.20 0.640 20.06 0.816 22.05 0.905
EnGAN [21] 61.01 114.35 17.48 0.650 18.23 0.617 16.57 0.734
RUAS [29] 0.83 0.003 18.23 0.720 18.37 0.723 16.55 0.652
FIDE [30] 28.51 8.62 18.27 0.665 16.85 0.678 15.20 0.612
KinD [18] 34.99 8.02 20.86 0.790 14.74 0.641 13.29 0.578
Restormer [12] 144.25 26.13 22.43 0.823 19.94 0.827 21.41 0.830
MIRNet [13] 785 31.76 24.14 0.830 20.02 0.820 21.94 0.876
SNR-Net [22] 26.35 4.01 24.61 0.842 21.48 0.849 24.14 0.928
Retinexformer [10] 15.57 1.61 25.16 0.845 22.80 0.840 25.67 0.930
  LYT-Net 3.49 0.045 24.13 0.844 22.93 0.840 23.33 0.905
Table 1: Quantitative results on LOL datasets and method complexity. Best results are in red, second best are in blue, while third best are in green.
@@ 3.4 Loss Function
$L_{\text{S}}=\begin{cases}0.5\cdot(y_{\text{true}}-y_{\text{pred}})^{2}&,\text{if }|y_{\text{true}}-y_{\text{pred}}|<1\\ |y_{\text{true}}-y_{\text{pred}}|-0.5&,\text{otherwise}\end{cases}$ (8)

### v4

@@ 1 Introduction
In this paper, we propose a novel transformer-based approach, characterized by its low-weight design, that achieves state-of-the-art (SOTA) results in LLIE while maintaining computational efficiency. In Fig. 1 we present a comparative analysis of performance over complexity between SOTA methods evaluated using the LOL dataset [2]. For visual clarity, in the short comparison, models such as Restormer[10] and MIRNet[11] are represented symbolically, accompanied by their detailed data.
@@ 3.3 Channel-wise Denoiser Block
FLOPS (G) Params (M) Original GTMean Original GTMean Original GTMean
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
  SID [23] 13.73 7.76 14.35 0.436 - - 13.24 0.442 - - 15.04 0.610 - -
3DLUT [24] 0.075 0.59 14.35 0.445 21.35 0.585 17.59 0.721 20.19 0.745 18.04 0.800 22.17 0.854
DeepUPE [25] 21.10 1.02 14.38 0.446 - - 13.27 0.452 - - 15.08 0.623 - -
DeepLPF [26] 5.86 1.77 15.28 0.473 - - 14.10 0.480 - - 16.02 0.587 - -
UFormer [27] 12.00 5.29 16.36 0.771 - - 18.82 0.771 - - 19.66 0.871 - -
RetinexNet [2] 587.47 0.84 16.77 0.560 18.92 0.427 15.47 0.567 18.32 0.447 17.13 0.798 19.09 0.774
Sparse [28] 53.26 2.33 17.20 0.640 - - 20.06 0.816 - - 22.05 0.905 - -
EnGAN [21] 61.01 114.35 17.48 0.650 20.00 0.691 18.23 0.617 - - 16.57 0.734 - -
RUAS [29] 0.83 0.003 18.23 0.720 18.65 0.518 18.37 0.723 19.06 0.510 16.55 0.652 16.58 0.719
FIDE [30] 28.51 8.62 18.27 0.665 - - 16.85 0.678 - - 15.20 0.612 - -
KinD [17] 34.99 8.02 20.86 0.790 20.86 0.802 14.74 0.641 17.54 0.669 13.29 0.578 16.25 0.591
Restormer [10] 144.25 26.13 22.43 0.823 26.68 0.853 19.94 0.827 26.12 0.853 21.41 0.830 25.43 0.859
MIRNet [11] 785 31.76 24.14 0.830 26.52 0.856 20.02 0.820 27.17 0.865 21.94 0.876 25.96 0.898
SNR-Net [22] 26.35 4.01 24.61 0.842 - - 21.48 0.849 - - 24.14 0.928 - -
Retinexformer [18] 15.57 1.61 25.16 0.845 27.14 0.850 22.80 0.840 27.69 0.856 25.67 0.930 28.99 0.939
  LYT-Net 3.49 0.045 22.38 0.826 27.23 0.853 21.83 0.849 27.80 0.873 23.78 0.921 29.38 0.939
Table 1: Quantitative results on LOL datasets and method complexity. Best results are in red, second best are in blue, while third best are in green. Original/GT Mean represent evaluations without/with gamma correction relative to the GT
@@ 3.4 Loss Function
$L_{\text{S}}=\begin{cases}F0.5\cdot(y_{\text{true}}-y_{\text{pred}})^{2}&,\text{if }|y_{\text{true}}-y_{\text{pred}}|<1\\ |y_{\text{true}}-y_{\text{pred}}|-0.5&,\text{otherwise}\end{cases}$ (8)